Jeff is 5 years older than Maria. Bob is 3 years older than twice Maria's age. Jeff, Maria, and Bob's ages added together equal 42. How old is Jeff?
Answer by ankor@dixie-net.com(22740)

ages represented by j, m and b, write an equation for each statement
:
Jeff is 5 years older than Maria.
j = m+5
Bob is 3 years older than twice Maria's age.
b = 2m+3
Jeff, Maria, and Bob's ages added together equal 42.
j + m + b = 42
Replace j and b using the 1st two equations
(m+5) + m + (2m+3) = 42
4m + 8 = 42
4m = 42 - 8
4m = 34
m = 34/4
m = 8.5 yrs is Marie
:
j = 8.5 + 5
j = 13.5 yrs is Jef
:
Usually get in integer solution to these, but this one is the exception.